Describe two laws of exponents and provide an example illustrating each law. Explain how to simplify your expression. How do the laws work with rational exponents? Provide a third expression to simplify that includes rational (fractional) exponents.

Answer by richwmiller(17219)

(x^2)^3=x^6
an exponent raised to a power is equal to the powers product of the exponents
x^2*x^3=x^5
to multiply two numbers raised to different powers add the exponents.
x^(1/2) is the same as getting the sqrt(x)
x^(1/2)=sqrt(x)